The density of ethanol is 0.789 g/ml at 20⁰C. Find the mass of a sample of ethanol that has a volume of 150.0 ml at this temperature

1
Expert's answer
2022-01-25T11:25:01-0500

D= M/V

0.789= M/150

M= 0.789×150

= 118.35g